In SATP to increase 1 mole gas volume to 40L, how much temperature should increase if pressure is half decreased---
1.389.43K
2.91.43°C
3.298K
4.116.43°C
1
Expert's answer
2019-01-18T05:46:52-0500

Find volume of 1 mole of a gas at SATP:

P= 101325 Pa

n= 1 mole

R= 8.314 m3Pa/(K*mole)

T = 298.15

PV=nRT

V= nRT/V = 1*8.314*298.15/101325 = 24.5 dm3

Use Combined Gas Law to determine final temparature:

P1V1/T1 = P2V2/T2

P2 = P1/2

P1*24.5 /298.15 = (P1/2) * 40 / T2

24.5/298.15 = 20/T2

T2 = 243.39 K

So, we can see that there is no increase in temperature as it is said in the task. There is decrese in temperature: T1 = 298.15 K and T2 = 243.39 K. The answer is: none of the above. Or there should be a mistake in the very task.